OpenStudy (anonymous):

What are the coordinates of the center and the radius of the circle given by the eqation? (x-1)^2 + (y-5)^2=16 Equation of a circle=(x-h)^2=(y-k)^2=r^2 Center: (h,k)

OpenStudy (campbell_st):

ok the general form is \[(x-h)^2 + (y - k)^2 = r^2\] any ideas of the values for h and k when you compare it to your equation?

OpenStudy (campbell_st):

and the 2nd part... the radius... in your question r^2 = 16 what is the value for r?